AAVSOnet news 2014-02-13
Recently, we've gone through all of the systems and indentified improvements to get file naming and data transfer consistent. That will make it simpler for the pipeline processing to handle incoming images. Notes on individual telescopes are given below.
BSM_Hamren has a focuser problem. Once solved, this system will be moved to another site - stay tuned!
BSM_Argentina is undergoing commissioning. A new V filter has been shipped to Jaime Garcia and should be installed in another week or so.
BSM_Berry is operational, now that the ST-8 camera has been repaired by SBIG. We are inserting an SA200 grating into the filter wheel in place of the clear filter for testing. In the spring, when the snow has melted from the roof, this system will be moved to another site - stay tuned!
BSM_NM is operational. 40 nights of data have been added to the epoch photometry database.
BSM_South is operational. All of its nights of data have been added to the epoch photometry database.
Coker30 is operational, but northern CA has finally gotten into a winter pattern and has been cloudy with rain for the past two weeks. Observing will resume as soon as it clears.
Bareket35 is undergoing commissioning. We have sent them new filters to replace their degraded originals.
SRO is fully operational.
TMO61 is fully operational, though there are some mechanical quirks that impact a few images, such as trailing or loss of telescope communication. These are being investigated. We will be inserting an SA200 grating into the filter wheel in another week or two for testing.
OC61 is fully operational. Mt. John staff are designing the instrument selector, so that the imager and spectrograph can be used with rapid changeover. We've sent them the guide camera for the spectrograph, and will be shipping the science camera next week.